During the current year, the liabilities of Williams Company decreased by $16,000 and the owners’ equity increased by $8,000.  The assets must have:

A.

decreased by $24,000

B.

None of the above

C.

increased by $24,000

D.

decreased by $8,000

E.

increased by  $8,000
Answer #1

Let beginning assets=100,000

Beginning liabilities=50,000

Beginning equity=50,000

Hence ending liabilities=(50,000-16000)=$34000

Ending equity=(50,000+8000)=$58000

Hence ending assets=(34000+58000)=$92000

Hence decrease in assets=(100,000-92000)=$8000

NOTE:

Total assets=Total liabilities+Total equity